Task #22814: 2.6 cavity with qubit (TTS1AES001)
Frequency detuning 2.6GHz
I am summarizing here the last finding on the 2.6GHz fundamental mode and detuning:
-- we see a change between the data from July 1st and July 24th. Looks like we are injecting more power.
-- I have checked every single component (cables, amplifier, splitter)
-- Last spectroscopy at lower power show instead a negative shift, see the normalized 2D plot of the low power sweep. However the detuning is very low about -40Hz
This is normalized power (ZOOM):
This is absolute value:
Roman showed about +100mK jump:
#2 Updated by Silvia Zorzetti over 1 year ago
- File 7376354398norm_mag.png 7376354398norm_mag.png added
- File 7376354398mag.png 7376354398mag.png added
- File 7376354398mag_lines.png 7376354398mag_lines.png added
From last spectroscopy there is a clear detuning, however this is negative (-50Hz).
I am running spectroscopy on the 4.7GHz mode to see if it has changed and will then resume the "qubit search".
#4 Updated by Silvia Zorzetti over 1 year ago
- File 7376356048mag.png 7376356048mag.png added
- File 7376356048norm_mag.png 7376356048norm_mag.png added
- File 7376356048mag_lines.png 7376356048mag_lines.png added
- File 7376356048freq_lines.png 7376356048freq_lines.png added
Here the finer scan of the 4.7GHz.
About 2.4kHz detuning.
This is the line cut at the bright state
Eric, I will shortly be in your office.